# Statistical features of quantum chaos using the Krylov operator complexity

### v1

@@ Abstract
Recently the Krylov operator complexity is proposed to evaluate the operator growth in quantum systems, and the variance of its Lanzcos coefficients is used as an important parameter for chaos. In this paper, we generate samples of random initial operators from given probability distribtions (GOE, GUE and the uniform distribution). For the Sinai billiard model, we study the statistical properties of the variance of Lanzcos coefficients in the associated Krylov space. Depending on whether the system is chaotic or not, the resulting distribution of the variance have different behaviors. In the nonchaotic case, the resulting distributions are almost overlapping together. In the chaotic case, they split into two well-separated groups. Besides, all the resulting distributions of the variance are the normal distribution, as long as the matrix size of the initial operator is large enough. This overlap-to-separation behavior, in the resulting distribution of the variance of Lanczos coefficients, may be a characteristics to distinguish chaotic and nonchaotic dynamics.
@@ 1 Introduction
The physics of complexity is an important topic that contributes to our understanding of both classical and quantum chaos, especially the growth of operator complexity. Recently the Krylov complexity Parker:2018yvk is proposed as an upper bound on the operator complexity, including the famous bound $\lambda_{L}\leq 2\pi k_{B}T/\hbar$ Maldacena:2015waa of the growth of the out-of-time-order correlation function (OTOC) 1969JETP28:1200L . The operator dynamics in the Krylov space resembles the hopping dynamics on the 1d chain Parker:2018yvk . From this, the connection between the dynamics of operator complexity and the Anderson localization is studied in Rabinovici:2021qqt , where the variance $\sigma^{2}$ of the Lanczos coefficients is connected to the late-time saturation of Krylov complexity.
@@
Then a statistical correlation between the Krylov complexity and the level statistics Dyson:1962JMP ; Gutzwiller:1971fy ; Berry:1977RSPSA ; Bohigas:1983er ; Blumel:1990PhRvL are firstly reported in spin systems Rabinovici:2021qqt ; Rabinovici:2022beu and billiard systems Hashimoto:2023swv . For the billiard system Hashimoto:2023swv , the variance of Lanczos coefficents is computed from the Krylov space of the momentum operator, and the level statistics are computed as usual. As the system Hamiltonian varies from nonchaotic to chaotic behavior, a sequence of data are generated for the level statistics and the variance of Lanczos coefficients. The statistical correlation coefficient between these two sequences of data are not zero. For the chaotic spin system Rabinovici:2022beu , the late-time saturation of Krylov complexity approaches the behavior of Random Matrix Theory (RMT) under certain circumstances. Also, the operator generating the Krylov space is randomly selected and three different operators are compared Rabinovici:2022beu . Recently in Craps:2024suj , the Krylov complexity is studied by choosing multiple initial operators, instead of one single operator.
@@
Along this line, it is a natural and straightforward question to ask what would happen if we do statistics on initial operators over a given probability distribution. In this paper, we study this possibility. We generate samples of initial operators from a probability distribution (GOE, GUE and the uniform distribution) and study the resulting distribution of the variance $\sigma^{2}$ of Lanzcos coefficients. Interestingly, we find that for the nonchaotic case, the resulting distributions of the variance $\sigma^{2}$ are nearly the same for different distributions of initial operators. And surprisingly, for the chaotic case, the resulting distributions of the variance $\sigma^{2}$ are obviously different. More intuitively, for different distributions of initial operators, the resulting plots of distributions are overlapped together in the nonchaotic case, and divide into two well-separated groups in the chaotic case.
@@
In detail, we consider a typical billiard system, the Sinai billiard, which originates from a quantization of classical chaos BERRY1981163 . The computation is done in the energy representation and operators become matrices. For the random initial operator $\mathcal{O}$ coming from the probability distribution GOE (GUE and the uniform distribution), the corresponding random matrix is sampled from the standard GOE (GUE and Hermitian matrices whose matrix elements follow the uniform distribution). Then the variance $\sigma^{2}$ of Lanczos coefficients is obtained in the associated Krylov space, when the system is chaotic or nonchaotic. By sampling the initial operator many times for each distribution, we can analyze the resulting statistical distribution of the variance $\sigma^{2}$. We find that the resulting distribution of $\sigma^{2}$ is almost always the normal distribution, independent of the initial operator distribution. As the system changes from nonchaotic to chaotic, the resulting normal distributions of $\sigma^{2}$ is computed and the logarithm of its statistical average $\langle\sigma^{2}\rangle$ and standard deviation are shown in Figure 5. In the nonchaotic case, the resulting distributions of $\sigma^{2}$ are almost overlapping together. In the chaotic case, they split into two groups well-separated from each other: one group consists of GOE, the uniform distribution of pure real numbers (URE) and of pure imaginary numbers (UIM), the other group is of GUE and the uniform distribution of complex numbers (UCP).
@@
The paper is organized as follows. In Section 2, we briefly review the Krylov operator space. In Section 3, we explain the sampling of the random initial operator. In Section 4, we give a qualitative explanation for the reason why the resulting distribution of the variance $\sigma^{2}$ is always the normal distribution. In Section 5, we show the detailed results of the model-the Sinai billiard. In Section 6, we discuss the physics of our results and conclude with a discussion of open questions.
@@ 2 Preliminary: Krylov operator space
An orthonormal basis, called the Krylov basis, can be constructed for $\mathcal{K}_{m}(\mathcal{L},\mathcal{O}(0))$. This construction also generates a sequence of important nubmers called the Lanczos coefficients, because the algorithm used is the Lanczos algorithm( or similar algorithms). The orthogonality is guaranteed by the inner product $(A|B)=\operatorname{tr}[A^{\dagger}B]$, which can be considered as a simplification of the Wightman inner product at infinite temperature. The Lanczos algorithm can be briefly stated as following:
@@
The Krylov basis is given by $\left\{\left|O_{0}\right),\left|O_{1}\right)\cdots\left|O_{n_{k}}\right)\right\}$ and the dimension of Krylov operator space is $n_{k}=dim(\mathcal{K}_{m})-1$. The sequence of positive numbers $\left\{b_{n}\right\}$ is called the Lanczos coefficient and is essentially the matrix element
@@
The Lanczos coefficient $\left\{b_{n}\right\}$ contains all the information about the evolution of the operator. To see this, firstly expand the operator $\mathcal{O}(t)$ in terms of the Krylov basis
@@
and then substitute it into the Heisenberg equation. This would give
@@
where the dot is the time derivative and the initial condition is $\varphi_{n}(0)=\delta_{n0}\|O\|$. The logarithmic variance $\sigma$ of the Lanczos coefficient $\left\{b_{n}\right\}$ is defined by
$\sigma^{2}\coloneqq\text{Var}(x_{i})=\left\langle x^{2}\right\rangle-\left\langle x\right\rangle^{2},x_{i}=\ln{\frac{b_{2i-1}}{b_{2i}}}.$ (2.7)
In Parker:2018yvk , the equation (2.6) is interpreted as the dynamincs on a 1d chain, with $\varphi_{n}(t)$ being the wave function and $b_{n}$ being hopping amplitude. Then the Krylov space can be associated with systems of spin chains. In this analog, the connection between the dynamics of complexity and the Anderson localization is studied in Rabinovici:2021qqt , and the importance of the variance of Lanczos coefficient is recognized.
3 Random initial Operators
In this paper, we choose samples of random initial operators from different probability distributions and study the resulting distribution of the variance $\sigma^{2}$ of the Lanczos coefficients. Specifically, we consider random initial operators ($\mathcal{O}_{GOE}$, $\mathcal{O}_{GUE}$ and $\mathcal{O}_{uni}$) from the Gaussian orthogonal ensemble (GOE), the Gaussian unitary ensemble (GUE) and the uniform distribution.
@@ 3 Random initial Operators
$P\left(\mathcal{O}^{(0)}_{11},\mathcal{O}^{(0)}_{12},\cdots,\mathcal{O}^{(0)}_{NN}\right)=(\frac{1}{2\pi})^{\frac{N}{2}}(\frac{1}{\pi})^{\frac{N^{2}-N}{2}}\exp\left[-\frac{1}{2}\text{Tr}(\mathcal{O}^{(0)2})\right]$ (3.2)
$P\left(\mathcal{O}^{(1)}_{11},\mathcal{O}^{(1)}_{12},\cdots,\mathcal{O}^{(1)}_{NN}\right)=(\frac{1}{2\pi})^{\frac{N}{2}}(\frac{1}{\pi})^{\frac{N^{2}-N}{2}}\exp\left[-\frac{1}{2}\text{Tr}(\mathcal{O}^{(1)2})\right]$ (3.3)
@@
For convenience, we introduce the distribution function $f_{\sigma^{2}}(\mathcal{O})$, whose value is the variance $\sigma^{2}$ of the Lanczos coefficients generated by the Krylov subspace $\mathcal{K}_{m}(\mathcal{L},\mathcal{O})$ of the initial operator $\mathcal{O}$
$f_{\sigma^{2}}(\mathcal{O})\coloneqq\text{the variance }\sigma^{2}\text{ of Lanczos coefficients of the Krylov subspace }\mathcal{K}_{m}(\mathcal{L},\mathcal{O}).$ (3.4)
Obviously, the resulting distribution function $f_{\sigma^{2}}$ depends on the initial distribution of the operator $\mathcal{O}$. Interestingly, when compare different initial probability distributions of $\mathcal{O}$, the behavor of the resulting distribution function $f_{\sigma^{2}}$ depends on whether the system is nonchaotic or chaotic: nearly complete overlapping $f_{\sigma^{2}}$ in the nonchaotic case, and two well-separated groups of $f_{\sigma^{2}}$ in the chaotic case. For all cases, the distribution function $f_{\sigma^{2}}$ is a normal distribution.
@@ 4 Statistical properties of Lanczos coefficients
Before showing detailed results, we give a qualitative analysis of the reason why the resulting distributions are normal distributions. This is related with the nature of the Krylov algorithm itself and can be understood using the central limit theorem of probability theory, so we explain it here in this section.
Firstly, we prove the following lemma by mathematical induction, which essentially says that the matrices involved in the Krylov algorithm depends only on the energy levels and the initial random matrix.
@@
From this lemma, we see that all Lanczos coefficients $b_{k}=b_{k}(O_{mn}^{(0)};E_{mn})$ depend only on the energy levels and the initial random operator $O_{mn}^{(0)}$. In practical computation, the energy levels are truncated to $\mathcal{N}_{max}$ and the random matrix $\mathcal{O}^{(0)}_{mn}$ is of size $\mathcal{N}_{max}\times\mathcal{N}_{max}$. If the value of $\mathcal{N}_{max}$ is large, we would have large-enough random numbers, so qualitatively the resulting distribution can be estimated by the central limit theorem. This is one aspect related to the large value of $\mathcal{N}_{max}$.

### v2

@@ Abstract
Recently the variance of Lanzcos coefficients in the Krylov space of an initial operator is recognized as an important indicator of quantum chaos. In this paper, we generate samples of random initial operators from given probability distribtions (GOE, GUE and the uniform distribution) and do statistics on the variance of Lanczos coefficients over these initial operators. Using the Sinai billiard with an integrability-breaking term, we propose two statistical quantities that have distinct behaviors as the system changes from nonchaotic to chaotic. One is the average correlation matrix $\langle x_{i}x_{j}\rangle$ of Lanczos coefficients, which exhibits different patterns in the nonchaotic and the chaotic regime. The other one is the resulting distribution of the variance of Lanzcos coefficients. In the nonchaotic case, the resulting distributions are almost overlapping together. In the chaotic case, they split into two well-separated groups. Furthermore, the resulting statistics of these two quantities are the Wishart distribution and the (rescaled) chi-square distribution respectively, which are independent of the distributions of initial operators and become the normal distribution in the case of large matrix size.
@@ 1 Introduction
The physics of complexity is important for understanding many different phenomena, ranging from condensed matter physics to black hole physics. Especially, the operator complexity captures the spread of an operator under time evolution and is an important indicator in the analysis of chaos. Recently the spread of an operator is explored in the Krylov space Parker:2018yvk , where the dynamics depends on the expansion coefficients of the operator in the Krylov basis and on the Lanczos coefficients generated in the process of constructing this Krylov basis. The Krylov complexity (K-complexity) Parker:2018yvk constructed from the expansion coefficients is proposed as an upper bound on the operator complexity, including the famous bound $\lambda_{L}\leq 2\pi k_{B}T/\hbar$ Maldacena:2015waa of the growth of the out-of-time-order correlation function (OTOC) 1969JETP28:1200L .
Interestingly, the operator dynamics in the Krylov space resembles the hopping dynamics on a 1d chain Parker:2018yvk , called the Krylov chain, with the hopping amplitudes being the Lanczos coefficients. Using this analog, the erratic behavior of Lanczos coefficients is analyzed Trigueros:2021rwj ; Rabinovici:2021qqt in terms of the Anderson localization. The variance $\sigma^{2}$ of Lanczos coefficients is introduced Rabinovici:2021qqt to explain the suppression of the late-time saturation value of K-complexity in the integrable XXZ spin chain. This suggests that the variance $\sigma^{2}$ of Lanczos coefficients may play an important role in distinguishing chaotic and nonchaotic behaviors.
Then it leads to an interesting connection between K-complexity and Random Matrix Theory (RMT). By adding an integrability-breaking term to the XXZ spin chain, the late-time saturation value of K-complexity is computed as the system changes from nonchaotic to chaotic behavior Rabinovici:2022beu . It approaches the behavior of similar values computed with Hamiltonian drawn from RMT ensembles Rabinovici:2022beu . Also, three different initial operators generating the Krylov space is randomly selected and compared there Rabinovici:2022beu . In Craps:2024suj , the K-complexity is also studied by choosing multiple initial operators.
@@
Furthermore, a statistical correlation between the variance $\sigma^{2}$ of Lanczos coefficients and the level statistics Dyson:1962JMP ; Gutzwiller:1971fy ; Berry:1977RSPSA ; Bohigas:1983er ; Blumel:1990zz is reported in the billiard systems Hashimoto:2023swv . As the system Hamiltonian varies from nonchaotic to chaotic, the variance $\sigma^{2}$ of Lanczos coefficients is computed for an initial momentum operator and the level statistics are computed as usual. The linear correlation coefficient between these two sequences of data are not zero.
@@
Along this line, an important question arises: what would happen for the Lanczos coefficients if we do statistics of initial operators over a given probability distribution. In this paper, we study this possibility by generating samples of initial operators from a given probability distribution (GOE, GUE and the uniform distribution) and analyzing the resulting Lanzcos coefficients. Interestingly, we find two statistical quantities that behave distinctively in the nonchaotic and the chaotic regime after averaging over the sample. One is the average correlation matrix $\langle x_{i}x_{j}\rangle$ of Lanczos coefficients. In the nonchaotic regime, it has a pattern that locally resembles the two-electron wave function of the Anderson localization Weinmann1995h2eOF , but this pattern disappears (or fading away by spreading out uniformly) in the chaotic regime. The other one is the resulting distribution of the variance $\sigma^{2}$. For samples of initial operators from different distributions, the resulting distributions of $\sigma^{2}$ are nearly the same in the nonchaotic case, but split into two well-separated groups in the chaotic case. Intuitively, the resulting plots of distributions of $\sigma^{2}$ are overlapped together in the nonchaotic case, and divide into two well-separated groups in the chaotic case. Furthermore, we find that the average correlation matrix $\langle x_{i}x_{j}\rangle$ satisfies the Wishart distribution and the resulting distribution of $\sigma^{2}$ satisfies a rescaled chi-square distribution. This is independent of the initial operator distributions and approaches the normal distribution as the size of the operator matrix increases.
@@
In detail, we consider the Sinai billiard, which is a typical billiard system originating from a quantization of classical chaos BERRY1981163 . The numerical computation is done in the energy representation and operators become matrices. Each time, a random initial operator $\mathcal{O}$ is sampled out by picking the corresponding random matrix from the following distributions: the GOE, the GUE, the uniform distribution of complex numbers (UCP), the uniform distribution of pure real numbers (URE) and of pure imaginary numbers (UIM). Then the Lanczos coefficients are computed in the Krylov space associated with this operator, which give the correlation matrix $x_{i}x_{j}$ and the variance $\sigma^{2}$ of Lanczos coefficients. Repeating this enough times, the average correlation matrix $\langle x_{i}x_{j}\rangle$ and the distribution of the variance $\sigma^{2}$ can be extracted for different samples of initial operators. Finally, these two statistical quantities are computed as the system changes from nonchaotic to chaotic. Independent of the initial distributions, the resulting statistics of $\langle x_{i}x_{j}\rangle$ and of $\sigma^{2}$ approach the normal distribution when the size of matrix becomes large, as discussed in Section 3. For the average correlation matrix $\langle x_{i}x_{j}\rangle$, its different patterns in the nonchaotic and the chaotic regime are shown in Figure 5. For the resulting normal distributions of $\sigma^{2}$, the logarithm of its statistical average $\langle\sigma^{2}\rangle$ and standard deviation are shown in Figure 8. In the nonchaotic case, the resulting distributions of $\sigma^{2}$ are almost overlapping together for all the five initial distributions. In the chaotic case, they split into two groups well-separated from each other: one group consists of GOE, URE and UIM, while the other group consists of GUE and UCP.
@@
The paper is organized as follows. In Section 2, we briefly review the relevant physics of the Krylov operator space and explain the sampling of random initial operators. In Section 3, we qualitatively explain the Wishart distribution of $\langle x_{i}x_{j}\rangle$ and the chi-square distribution of $\sigma^{2}$, which approach the normal distribution when the size of matrix is large enough. In Section 4, we show the detailed results for the Sinai billiard. In Section 5, we discuss the physics of our results and conclude with a discussion of open questions.
@@ 2 Preliminary: Krylov operator space
An orthonormal basis, called the Krylov basis, can be constructed for $\mathcal{K}_{m}(\mathcal{L},\mathcal{O}(0))$. This construction also generates a sequence of important nubmers called the Lanczos coefficients, because the algorithm used is the Lanczos algorithm( or similar algorithms). The orthogonality is guaranteed by the inner product $(A|B)=\operatorname{tr}[A^{\dagger}B]$, which can be considered as a simplification of the Wightman inner product at infinite temperature. The Lanczos algorithm can be briefly stated as follows:
@@
The Krylov basis is given by $\left\{\left|O_{0}\right),\left|O_{1}\right)\cdots\left|O_{n_{k}-1}\right)\right\}$ and the dimension of Krylov operator space is $n_{k}=dim(\mathcal{K}_{m})$. The sequence of positive numbers $\left\{b_{n}\right\}$ is called the Lanczos coefficient and is essentially the matrix element of the Liouvillian operator
@@
The Lanczos coefficient $\left\{b_{n}\right\}$ is essential for the evolution of the operator $\mathcal{O}(t)$. To see this, firstly expand the operator $\mathcal{O}(t)$ in terms of the Krylov basis
@@
and then substitute it into the Heisenberg equation, which leads to the following equation
@@
with the dot being the time derivative and the initial condition being $\varphi_{n}(0)=\delta_{n0}\|O\|$. This equation suggests that the spreading behavior of an operator in the Krylov space resembles the dynamics of wave functions on a 1d chain, called the Krylov chain Parker:2018yvk , where $\varphi_{n}(t)$ is the wave function and $b_{n}$ is the hopping amplitude. This makes it possible to connect the operator complexity in the Krylov space with condensed matter physics on spin chains. The tridiagonal form of the Liouvillian operator is another feature of the Krylov chain. On this Krylov chain, the time-dependent average position acts naturally as an indicator of the spreading behavior and is defined as the K-complexity
$C_{K}(t)=\sum_{n=0}^{n_{k}-1}n\left|\phi_{n}(t)\right|^{2}.$ (2.7)
The K-complexity has been playing an important role in the analysis of quantum chaos, see the review Nandy:2024htc for related works.
Obviously, results in the Krylov space will vary when the initial operator $\mathcal{O}$ changes. One way to deal with this dependence on initial operators is to do statistics and this is what we do in this paper. We choose samples of random initial operators from different probability distributions and study the resulting statistics of the Lanczos coefficients. Specifically, we consider random initial operators ($\mathcal{O}_{GOE}$, $\mathcal{O}_{GUE}$ and $\mathcal{O}_{uni}$) from the Gaussian orthogonal ensemble (GOE), the Gaussian unitary ensemble (GUE) and the uniform distribution.
@@
$P\left(\mathcal{O}^{(0)}_{11},\mathcal{O}^{(0)}_{12},\cdots,\mathcal{O}^{(0)}_{NN}\right)=(\frac{1}{2\pi})^{\frac{N}{2}}(\frac{1}{\pi})^{\frac{N^{2}-N}{2}}\exp\left[-\frac{1}{2}\text{Tr}(\mathcal{O}^{(0)2})\right],$
$P\left(\mathcal{O}^{(1)}_{11},\mathcal{O}^{(1)}_{12},\cdots,\mathcal{O}^{(1)}_{NN}\right)=(\frac{1}{2\pi})^{\frac{N}{2}}(\frac{1}{\pi})^{\frac{N^{2}-N}{2}}\exp\left[-\frac{1}{2}\text{Tr}(\mathcal{O}^{(1)2})\right],$ (2.9)
@@
$\mathcal{O}_{mn}^{(0,1)}\sim U(-\sqrt{3},\sqrt{3})$ (2.10)
@@ 3 Statistical properties of Lanczos coefficients
The perspective of the Krylov chain (2.6) provides an efficient way to analyze the spreading dynamics of an operator by connecting to the physics of quantum many-body systems. Firstly in Dymarsky:2019elm , the delocalization on the chain is proposed as the reason for the chaotic behavior of an operator on the many-body systems. Then in Rabinovici:2021qqt , the localization on the Krylov chain is proposed as the reason for the nonchaotic behavior of an operator on the XXZ spin chain.
Here, the Lanczos coefficients plays an an essential role because it is the hopping amplitude on the Krylov chain. The sequence of Lanczos coefficients is more erratic in the nonchaotic case and less erratic in the chaotic case. This behavior is explored Rabinovici:2021qqt ; Trigueros:2021rwj on the Krylov chain as a phenomenon of the Anderson localization. The energy eigenstate $\mid\omega)$ on the Krylov chain can be decomposed in the Krylov basis as
$\mid\omega)=\sum_{n=0}^{n_{k}-1}\psi_{n}\mid\mathcal{O}_{n})$ (3.1)
There is a zero frequency state whose solution is given by Rabinovici:2021qqt ; Trigueros:2021rwj
$\frac{\psi_{2n}}{\psi_{0}}=(-1)^{n}\prod_{i=1}^{n}\frac{b_{2i-1}}{b_{2i}},\quad n=0,\ldots,\frac{n_{k}-1}{2}.$ (3.2)
The extent of localization can be extracted from this zero frequency state and the disordered cofficients $b_{n}$ fleishman1977fluctuations . The localization length is inversely proportional to the variance of the Lanczos coefficients Rabinovici:2021qqt , defined by
$\sigma^{2}\coloneqq\text{Var}(x_{i}),\,x_{i}=\ln{\left|\frac{b_{2i-1}}{b_{2i}}\right|}.$ (3.3)
Intuitively, the variance $\sigma^{2}$ measures the magnitude of the erratic behavior of $b_{n}$. A smaller variance $\sigma^{2}$ means less erratic behavior of Lanczos coefficients and this gives a larger localization length, so the Krylov chain is more delocalized and this happens when the system is more chaotic. Results in Rabinovici:2021qqt show that the variance $\sigma^{2}$ is indeed smaller in the chaotic case.
[…]
To be specific, We focus on the statistics of the more interesting correlation matrix  [^2] In statistics, this is the scatter matrix to estimate the covariance matrix of a distribution. Since we are not doing pure mathematics here, we prefer to use the more physical term ’correlation matrix’ Eynard:2015aea . $x_{i}x_{j}$, and the resulting distribution $f_{\sigma^{2}}$ of the variance $\sigma^{2}$ across a sample. Naively, the resulting statistics of $x_{i}x_{j}$ and $f_{\sigma^{2}}$ depend on the distributions of the initial operator $\mathcal{O}$. However, we find that their statistics are the Wishart distribution and the rescaled chi-square distribution respectively, independent of the distribution of initial operators. As the matrix size of the operator increases, they approach the normal distribution, at which time the behavior of $f_{\sigma^{2}}$ depends only on whether the system is nonchaotic or chaotic: nearly complete overlapping $f_{\sigma^{2}}$ in the nonchaotic case, and two well-separated groups of $f_{\sigma^{2}}$ in the chaotic case.
Before showing the explicit results, we qualitatively analyze the statistical propertities of Lanczos coefficients in the remaining of this section. We firstly explain the statistics in a purely intuitive manner. Then we explain the statistics using the Wishart distribution wishart1928generalised .
3.1 Intuition on the normal distribution
Firstly we give an intuitive explanation of the reason why the resulting distributions are normal distributions when the size of matrix is large. This is related with the nature of the Krylov algorithm itself and can be understood using the central limit theorem of probability theory.
Intuitively, the matrices involved in the Krylov algorithm depend only on the energy levels and the initial random matrix. This can be seen by the following lemma proved by mathematical induction.
@@ 3.1 Intuition on the normal distribution
From this lemma, we see that all Lanczos coefficients $b_{k}=b_{k}(O_{mn}^{(0)};E_{mn})$ depend only on the energy levels and the initial random matrix $O_{mn}^{(0)}$. In practical computation, the energy levels are truncated to $\mathcal{N}_{max}$ and the random matrix $\mathcal{O}^{(0)}_{mn}$ is of size $\mathcal{N}_{max}\times\mathcal{N}_{max}$. If the value of $\mathcal{N}_{max}$ is large, there would be tremendous random numbers. Intuitively, we would expect that the central limit theorem comes into play when we have large-enough random numbers, and the resulting distribution can be estimated by the normal distribution. On the other hand, we can intuitively expect that most information of the system is preserved with a large enough $\mathcal{N}_{max}$. We know that a chaotic system resembles the behavior of RMT from level statistics. Then a more accurate approximation to the system means more randomness are preserved in this case. So things are random enough such that we can use the central limit theorem to expect the normal distribution.
3.2 Distribution of the average correlation matrix $\langle x_{i}x_{j}\rangle$
Now we qualitatively analyze the statistics of the average correlation matrix $\langle x_{i}x_{j}\rangle$. Suppose $K$ continuous Lanczos coefficients are selected from the Krylov space of a given initial operator, with $K$ being the algorithmic steps. We shall take the approximation that the sequence $\vec{X}=(x_{1},\ldots,x_{K})$ is a $K$-component random vector. This is easily understood in the Krylov chain, where by analog with the Anderson localization the Lanczos coefficients are disordered. Without the Krylov chain, it is indeed harder to accept this. Naively looking at the Lanczos algorithm, the Lanczos coefficients seem to be already predetermined by the initial operator, like an initial value problem of ODE. However, we can intuitively get the randomness from Lemma 1. The initial operator is already a random matrix and all Lanczos coefficients are generated from a complicated pipeline of arithmetic of energy levels with the random matrix. And the energy levels themselves are closely related with RMT in the case of chaos. So we can still accept the assumption that $\vec{X}$ is a random vector even without the help of the Krylov chain. After all, studies in the literature have shown the erratic behavior of Lanczos coefficients in various kinds of systems, and this assumption indeed leads to a successful interpretation of the results in this paper.
After accepting the randomness of $x_{i}$, we go one step further by assuming that the $K$-component random vector $\vec{X}=(x_{1},\ldots,x_{K})$ is from a multivariate normal distribution $N_{K}(0,\Sigma)$ with zero mean and a covariance matrix $\Sigma$ of size $K$. The zero mean is easily understood again in the Krylov chain by the Anderson localization, where the disordered $b_{n}$ imply that $x_{i}$ are fluctuating around zero. The property of being normal can be intuitively understood from the analysis in the previous subsection 3.1. Previous numerical results also show this property Trigueros:2021rwj ; Rabinovici:2022beu .
With this assumption, the average correlation matrix $\langle x_{i}x_{j}\rangle$ is exactly the common estimator used in RMT Eynard:2015aea . For a sample of $N$ random initial operators, we get $N$ random vectors that form a matrix of size $N\times K$
$X=\left(\vec{X}_{1},\vec{X}_{2},\cdots,\vec{X}_{N}\right)^{\mathrm{T}},\quad\vec{X}_{i}=\left(x^{(i)}_{1},x^{(i)}_{2},\cdots,x^{(i)}_{K}\right),\quad x^{(i)}_{j}\in\mathbb{R},$ (3.12)
where the $\vec{X}_{i}$ are independent, identically distributed random vectors from $N_{K}(0,\Sigma)$. Then the sample covariance matrix $M$ is a symmetric square matrix of size $K$ defined as
$M=X^{\mathrm{T}}X=\sum_{i=1}^{N}\vec{X}^{\mathrm{T}}_{i}\vec{X}_{i}\sim W_{K}(\Sigma,N),$ (3.13)
which has the probability distribution $W_{K}(\Sigma,N)$ known as the Wishart distribution. The empirical sample covariance matrix $M$ is the most commonly used estimator for the covariance matrix $\Sigma$. If $\Sigma$ is given, the random vector $\vec{X}_{i}$ has the distribution
$N_{K}(0,\Sigma)\left(\vec{X}_{i}\right)=\frac{e^{-\frac{1}{2}\operatorname{Tr}\vec{X}_{i}\Sigma^{-1}\vec{X}_{i}^{\mathrm{T}}}}{(2\pi)^{K/2}\sqrt{\operatorname{det}\Sigma}}\prod_{j=1}^{K}dx^{(i)}_{j}$ (3.14)
the Wishart distribution is defined as
$P(M)=\frac{1}{2^{NK/2}(\operatorname{det}\Sigma)^{N/2}\Gamma_{K}\left(\frac{N}{2}\right)}(\operatorname{det}M)^{\frac{N-K-1}{2}}e^{-\frac{1}{2}\operatorname{Tr}\Sigma^{-1}M},$ (3.15)
where $\Gamma_{K}$ is a multivariate generalization of the Gamma function
$\Gamma_{K}\left(\frac{N}{2}\right)=\pi^{K(K-1)/4}\prod_{j=1}^{K}\Gamma\left(\frac{N}{2}-\frac{j-1}{2}\right).$ (3.16)
Note that our average correlation matrix $\langle x_{i}x_{j}\rangle$ is exactly the sample covariance matrix $M$, because $\langle x_{i}x_{j}\rangle=M_{ij}/N$. So it can indeed capture the statistics behind the ensemble average and we can use properties of the Wishart distribution to understand its behavior. The Wishart distribution can be viewed as a multi-variate generalization of the chi-squared distribution as follows: if $K=\Sigma=1$, it reduces to the chi-squared distribution with $N$ degrees of freedom. However, this is exactly the case in this paper. The distributions of initial operators we used (2.8)-(2.10) are univariate, whose $\Sigma$ is the identity matrix. Then it is naturally to expect that the resulting $K$-component random vector $\vec{X}$ is also univariate. So we can refine the previous assumption by expecting that $\vec{X}$ is from a univariate normal distribution $N(0,\sigma_{x})$. In this way, the Wishart distribution satisfied by $\langle x_{i}x_{j}\rangle$ reduces to the chi-square distribution of $\langle x_{i}^{2}\rangle$. In Figure 1(a), we show the histogram of $\langle x_{i}^{2}\rangle$ for the case of GUE when the system is chaotic with $\mathcal{N}_{max}=15$. It is indeed the chi-square distribution. As the size of matrix increases, the distribution would become normal by the central limit theorem. In Figure 1(b), the corresponding histogram of $\langle x_{i}x_{j}\rangle$ with $\mathcal{N}_{max}=50$ is plotted and it is indeed the normal distribution. Similarly properties are there when the system is nonchaotic.
(a) Histogram of $\langle x_{i}^{2}\rangle$, $\mathcal{N}_{max}=15$.
(b) Histogram of $\langle x_{i}x_{j}\rangle$, $\mathcal{N}_{max}=50$.
Figure 1: Distribution of $\left\langle x_{i}x_{j}\right\rangle$ in the chaotic case.
3.3 Distribution $f_{\sigma^{2}}$ of the variance $\sigma^{2}$
Now we have a qualitative understanding of the statistics of the average correlation matrix $\left\langle x_{i}x_{j}\right\rangle$ from the aspect of the Wishart distribution. What about the statistics of the resulting distribution $f_{\sigma^{2}}$ of the variance? We expect it to follow the same distribution as the average correlation matrix.